-4/6z + 2=30, I need to help my son write this out and be able to explain the steps to him. thanks in advance for the help
Answer by checkley71(8403)

-4/6Z+2=30
-4/6Z=30-2
-4/6Z=28 CROSS MULTIPLYING WE GET
-4Z=28*6
-4Z=168
Z=168/-4
Z=-42
PROOF
(-4/6)-42+2=30
(-4*-42)/6+2=30
168/6+2=30
28+2=30
30=30